OpenStudy (anonymous):

line through point (4,3) that is perpendicular to the line y=4x-2

OpenStudy (mathstudent55):

What do you know about the slopes of perpendicular lines?

OpenStudy (anonymous):

the line y=4x-2 the slope here is m=4 for perpendicular line its slope is m2=-1/m1 m2=-1/4 now use the point (4,3) and m=-1/4 to get the equation of the line using y-y1=m(x-x1)